# An Analytic Characterization of the Limb Asymmetry - Transit Time Degeneracy

### v1

@@ Abstract
Atmospheres are not spatially homogeneous. This is particularly true for hot, tidally locked exoplanets with large day-to-night temperature variations, which can yield significant differences between the morning and evening terminators – known as limb asymmetry. Current transit observations with the James Webb Space Telescope (JWST) are precise enough to disentangle the separate contributions of these morning and evening limbs to the overall transmission spectrum in certain circumstances. However, the signature of limb asymmetry in a transit light curve is highly degenerate with uncertainty in the planet’s time of conjunction. This raises the question of how precisely transit times must be measured to enable accurate studies of limb asymmetry, particularly with JWST. Although this degeneracy has been discussed in the literature, a general description of it has not been presented. In this work, we show how this degeneracy results from apparent changes in the transit contact times when the planetary disk has asymmetric limb sizes. We derive a general formula relating the magnitude of limb asymmetry to the amount which it would cause the apparent time of conjunction to vary, which can reach tens of seconds. Comparing our formula to simulated observations, we find that numerical fitting techniques add additional bias to the measured time, of generally less than a second, resulting from the occultation geometry. We also derive an analytical formula for this extra numerical bias. These formulae can be applied to planning new observations or interpreting literature measurements, and we show examples for commonly studied exoplanets.
facilities: Exoplanet Archivesoftware: batman (Kreidberg, 2015), catwoman (Espinoza & Jones, 2021; Jones & Espinoza, 2022), emcee (Foreman-Mackey et al., 2013), matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), PandExo (Batalha et al., 2017), SciPy (Virtanen et al., 2020)
@@ 1 Introduction
Atmospheres are not perfectly uniform. Observational surveys of brown dwarfs and planetary mass objects have found robust evidence that their atmospheres are heterogeneous and variable (e.g., Buenzli et al., 2014; Metchev et al., 2015; Apai et al., 2017). For transiting exoplanets specifically, there is abundant evidence for day-to-night variation in atmospheric temperature and cloudiness (e.g., Beatty et al., 2019; May et al., 2022, among many others) driven by the contrast in irradiation received by the day and night hemispheres. 3D atmospheric circulation models also predict that this day-to-night contrast will drive differences between the morning and evening terminator regions (e.g., Kataria et al., 2016; Line & Parmentier, 2016; Powell et al., 2019), which is known as “limb asymmetry” since these terminators form the limbs of the planetary disk when seen in transit.
@@
The James Webb Space Telescope (JWST) has proven capable of directly detecting limb asymmetry at low spectral resolution as well. The observational signature of limb asymmetry in JWST transit observations are apparent time of conjunction variations with wavelength (Feinstein et al., 2022; Rustamkulov et al., 2022) and changes in the chromatic transit depth between ingress and egress (Murphy et al., 2024; Espinoza et al., 2024; Delisle & et al., 2024, in review). One advantage of JWST is that it captures the transmission spectra, both overall and of each limb’s contribution, over a much wider wavelength range than is generally possible in high resolution spectroscopy from the ground. This enables more detailed retrievals of the atmospheric temperature-pressure structure, molecular abundances, and aerosol properties in each spatial region using JWST. Extracting these separate contributions is vital to studying and understanding the circulation of heat and molecules within atmospheres. Also, previous analyses have shown that if such differences are ignored, the properties retrieved from observed data can be biased significantly from the truth (Feng et al., 2016; Caldas et al., 2019; Taylor et al., 2020).
@@ 2.1 Defining the problem
Consider an example planet with asymmetric limbs, where each limb is represented by a semi-circle joined to the other along its flat edge (as in Figure 1, inspired by Espinoza & Jones, 2021; Jones & Espinoza, 2022). The evening limb has radius $R_{\mathrm{evening}}$ and the morning limb radius differs by some amount $\Delta R$, so that $R_{\mathrm{morning}}=R_{\mathrm{evening}}+\Delta R$. This $\Delta R$ can be positive or negative depending on which limb is larger, and herein we will assume $\Delta R$ is positive, so that the example planet’s morning limb is larger. While this construction simplifies the geometry, we note it is not a strictly realistic model as it creates a discontinuity in the atmosphere at the planet’s poles which grows with $\Delta R$. However, this simplified construction is likely adequate given current data quality, which we discuss later.
An observer seeking to study this planet’s limb asymmetry must measure both $R_{\mathrm{evening}}$ and $R_{\mathrm{morning}}$ [^1] In practice, it is really the corresponding planet-star radius ratios $R_{\mathrm{evening}}/R_{\star}$ and $R_{\mathrm{morning}}/R_{\star}$ that are measured from a transit light curve. as a function of wavelength, thus constructing separate transmission spectra for each limb. We assume that previous observers of this planet were not aware of – or not capable of – measuring limb asymmetry on the planet, and fit their data assuming the planet had uniform limbs. The planetary radius that these analyses will have inferred is the radius of a uniform circle with the same occulting area – the same transit depth – as the asymmetric-limb disk, which we call the “effective radius” $R_{\mathrm{eff}}$, calculated as
@@
Figure 1: Exaggerated schematic illustrating the difference in the transit light curves of an asymmetric-limb planet and a uniform-limb planet. Limb darkening is omitted for visual simplicity. Both planets share the same time of conjunction. Due to the difference in their leading and trailing limb radii though, the transit contact times of each planet differ. An observer that measures the light curve of the asymmetric limb planet, which is shown in blue, and fits their data using a uniform-limb transit model would infer a biased time of conjunction because these changes in the contact times effectively shift the transit earlier or later in time. This is the origin of the degeneracy between limb asymmetry and variations or uncertainty in a planet’s time of conjunction.
To best intuit the degeneracy between $\Delta R$ and $\Delta t_{c}$, it is useful to look at the light curve of the asymmetric-limb planet compared to its uniform-limb counterpart. Figure 1 shows an exaggerated example of each at the bottom, which are aligned in time with the sequence of disk schematics at the top. In this example, both planets have the exact same true time of conjunction and orbital speed, and are moving from left to right across the stellar disk. The asymmetric-limb planet will begin to transit before the uniform-limb planet because its leading limb, which is $R_{\mathrm{morning}}$, is larger than $R_{\mathrm{eff}}$. In other words, because it has a larger limb, the asymmetric-limb disk makes first contact (T1) with the stellar disk before the uniform-limb disk does. Similarly, since the asymmetric-limb disks’ trailing edge, $R_{\mathrm{evening}}$, is smaller than $R_{\mathrm{eff}}$, it is still “ahead” of the trailing edge of the uniform-limb disk so it will also contact the stellar disk first during the end of ingress (T2). The exact same sequence of events occurs during transit egress for the third (T3) and fourth (T4) contact points. Also note that the shape of the ingress and egress in each light curve is slightly different due to the different shape of each disk, but they both reach the same depth during full transit since both disks occult the same total area.
@@ 2.2 Times of first and last contact
We assume that each planet rotates about a single axis which is oriented parallel to its orbital axis and lies in the plane of the page. The following derivation will hold reasonably well even if there is small rotation-orbital axis misalignment, though such misalignment is unlikely for short-period exoplanets due to tidal interactions with the host star (e.g. Hut, 1981; Correia & Laskar, 2010). As shown in Figure 1, we construct each planetary disk as two semi-circles whose straight edges are joined along the planet’s rotation axis. Here, we explicitly define the “center” of each planet as the point where the rotation axis meets the two lines which bisect the arcs of each semi-circle, as indicated in Figure 1. Under this construction, these bisectors are parallel to the planets’ transit trajectories. To aid our derivation, we will work in a cartesian coordinate system as is also shown in Figure 1. The x-axis is aligned along the planet’s transit chord, so that the center of the planet is always moving directly along the x-axis.
The point of conjunction is the position of the center of each planet at its time of conjunction, defined as when the center of the planet is closest to the center of the stellar disk. We define this point of conjunction to be the origin of our coordinate system, so that the y-axis passes through this point and is orthogonal to both planets’ transit trajectories. Note that in Figure 1, we assume both planets transit directly along the stellar equator but this need not be true in general. Our coordinate system is invariant to general changes in the planets’ orbital orientations as the x-axis can be translated vertically along the stellar disk and the y-axis rotated such that the x-axis is always horizontal.
@@
Figure 2: Projected geometry of a planet with asymmetric limbs at transit first contact. A line can be drawn from the center of the star to the center of the planet which passes through the point of first contact. This line defines the hypotenuse of a triangle which we use to derive how the time of first contact changes as the leading limb radius changes in Equation 2.2.
When the asymmetric-limb planet makes first contact, assuming it has a larger leading limb, the uniform-limb planet will still need to travel some extra projected distance in the x-direction before it too makes first contact. Let us hold the asymmetric-limb planet fixed at first contact, and let the uniform-limb planet move forward in time until it too makes first contact. Then, we solve for the difference in these times via this extra projected distance $\Delta x$ that the uniform-limb planet travels. When the planets are transiting along the stellar equator, as is shown in Figure 1, this is simply $\Delta x=R_{\mathrm{morning}}-R_{\mathrm{eff}}\approx\Delta R/2$. For non-zero impact parameters, estimating $\Delta x$ is more involved. Figure 2 shows the projected geometry at the point of first contact for an asymmetric-limb planet with non-zero impact parameter. We can draw a triangle with a hypotenuse from the center of the stellar disk, through the point of first contact, to the center of the planetary disk. The length of this hypotenuse is $R_{\star}+R_{\mathrm{morning}}$ and it makes an angle $\theta_{1}$ with respect to the stellar equator. The projected vertical distance between the center of the planet and the stellar equator is $bR_{\star}$, where $b$ is the impact parameter. Let us define the position of the center of the asymmetric-limb planet when its leading limb makes first contact as $x_{a}$, and likewise for the uniform-limb planet as $x_{u}$. The extra distance travelled by the uniform-limb planet is {dmath} Δx = x_a - x_u
@@
Here, $R_{p,i}$ can be $R_{\mathrm{morning}}$ or $R_{\mathrm{eff}}$. We can convert Equation 3 to the cosine of the angle,
@@
The same derivation can be done for the time of fourth contact by swapping the morning limb radius for the evening limb radius. In our construction, the uniform-limb planet will also make fourth contact after the asymmetric-limb planet, now because the uniform-limb radius is larger than the evening-limb radius. Changing out terms, we then have {dmath} ΔT_4 = P2 πa ( (R_⋆+ R_eff )^2 - (b R_⋆)^2 - (R_⋆+ R_evening )^2 - (b R_⋆)^2 ). By construction, $\Delta T_{1}$ and $\Delta T_{4}$ will always have the same sign. This is because the equivalent-depth uniform radius will always be between the two limb radii, so T1 and T4 will always be earlier or later together. If the asymmetric-limb planet’s morning limb is larger than its evening limb, as we have been using in our constructed example, then $\Delta T_{1}$ and $\Delta T_{4}$ will be negative. If the evening limb is larger, they will be positive. Then, because these contact times are equidistant about the time of conjunction for the uniform-limb planet, {dmath} Δt_c = 12 ( ΔT_1 + ΔT_4 ).
@@ 3.1 Setup
We first simulated a “perfect” transit observation of the exoplanet WASP-39b in order to directly compare to our analytical formulae. We generated a noiseless light curve with data points at a cadence of 0.5 seconds. For this “perfect” observation, we finely sampled $N$ in twenty evenly-spaced steps between 0 and 5.
@@
Parameter [units] WASP-80 b WASP-96 b WASP-39 b WASP-107 b
P [days] 3.06785271(19)aafootnotemark: 3.4252565(8)aafootnotemark: 4.0552941(34)fffootnotemark: 5.721488(3)aafootnotemark:
@@
$i$ [degrees] 89.02 $\pm$ 0.11bbfootnotemark: 85.6 $\pm$ 0.2ccfootnotemark: 87.32 $\pm$ 0.17fffootnotemark: 89.7 $\pm$ 0.2jjfootnotemark:
@@
R [R] 0.605 $\pm$ 0.048**footnotemark: 1.15 $\pm$ 0.03**footnotemark: 0.92 $\pm$ 0.03**footnotemark: 0.73 $\pm$ 0.15**footnotemark:
@@
Teq [K] 825 $\pm$ 19bbfootnotemark: 1285 $\pm$ 40ccfootnotemark: 1166 $\pm$ 14fffootnotemark: 770 $\pm$ 60jjfootnotemark:
T14 [hours] 2.131 $\pm$ 0.003bbfootnotemark: 2.4264 $\pm$ 0.0264ccfootnotemark: 2.8032 $\pm$ 0.0192iifootnotemark: 2.753 $\pm$ 0.007jjfootnotemark:
H [km] 226 $\pm$ 44 669 $\pm$ 97 1042 $\pm$ 171 1221 $\pm$ 174
Cadence [s] 14.8 72.0 79.5 20.2
$\sigma$ [ppm] 200 150 200 200
@@
We then simulated more realistic transit observations of each planet using the optimal integration cadences computed using PandExo (Batalha et al., 2017) for a JWST/NIRCam F322W2 observation, and light curve scatters of 200 ppm. For these simulations, we used a coarser sampling for $N$ of ten steps between 0 and 5. The cadences are given in Table 1 and ranged from $\sim$15 - 80 seconds, consistent with those used by real observations of these targets (Bell et al., 2023; Ahrer et al., 2023; Murphy et al., 2024). This light curve scatter is slightly optimistic for JWST/NIRCam F322W2, as real observations of these targets have achieved $\sim$230-500 ppm scatter (Bell et al., 2023; Carter et al., 2024; Murphy et al., 2024), but it has been achieved by other JWST instruments like NIRISS SOSS (Radica et al., 2023). For WASP-96b, its nominal cadence of 144 s and 200 ppm light curve scatter led to large noise in our simulation results. We therefore used a 2$\times$ faster cadence of 72 s and slightly smaller scatter of 150 ppm just for WASP-96b, which did not change the underlying result but made it easier to visually present and compare to our analytical results. In all cases, we assumed having an equal time out-of-transit baseline to the published transit duration.
@@
After generating the simulated observed data, we fit these data using a uniform-limb transit model generated by the transit modelling package batman (Kreidberg, 2015), also with uniform limb darkening. We left all parameters fixed to their true values except for the time of conjunction, which was allowed to vary freely. We set the planetary radius to the value which would have the same planetary disk occulting area as the corresponding asymmetric-limb disk, using Equation 1. In the “perfect” noiseless case, we simply fit the data through a least squares regression. Then in the other cases, we fit the data using the Markov Chain Monte Carlo method using emcee (Foreman-Mackey et al., 2013), sampling just the time of conjunction for 5,500 steps which was more than sufficient for each run’s sampling to converge.
@@
We chose to use uniform limb darkening when generating the reference light curves, meaning the stellar disks are completely uniform in brightness. This choice was partially motivated by computational efficiency, as the overhead initializations of catwoman and batman are significantly slower when limb darkening must be computed. It was also motivated by our desire to treat each planet uniformly, as different host stars exhibit different limb darkening that may yield different results depending on the assumed cadence or noise of the light curve. We did test various limb darkening laws, including quadratic and logarithmic, and various sets of limb darkening coefficient values. We found that the precise choice of limb darkening did not change our result so long as it was well known, because we set the limb darkening when generating the reference light curve and kept it fixed for all later steps, thus assuming perfect knowledge of limb darkening during the light curve fitting. Uncertainty in the limb darkening coefficients is of concern to real observations, but is beyond the scope of our analysis and can be mitigated using stellar models in practice.

### v2

@@ Abstract
Atmospheres are not spatially homogeneous. This is particularly true for hot, tidally locked exoplanets with large day-to-night temperature variations, which can yield significant differences between the morning and evening terminators – known as limb asymmetry. Current transit observations with the James Webb Space Telescope (JWST) are precise enough to disentangle the separate contributions of these morning and evening limbs to the overall transmission spectrum in certain circumstances. However, the signature of limb asymmetry in a transit light curve is highly degenerate with uncertainty in the planet’s time of conjunction. This raises the question of how precisely transit times must be measured to enable accurate studies of limb asymmetry, in particular with JWST. Although this degeneracy has been discussed in the literature, a general description of it has not been presented. In this work, we show how this degeneracy results from apparent changes in the transit contact times when the planetary disk has asymmetric limb sizes. We derive a general formula relating the magnitude of limb asymmetry to the amount by which it would cause the apparent time of conjunction to vary, which can reach tens of seconds. Comparing our formula to simulated observations, we find that numerical fitting techniques add additional bias to the measured time, of generally less than a second, resulting from the occultation geometry. We also derive an analytical formula for this extra numerical bias. These formulae can be applied to planning new observations or interpreting literature measurements, and we show examples for commonly studied exoplanets.
journal: The Astrophysical Journalfacilities: Exoplanet Archivesoftware: batman (Kreidberg, 2015), catwoman (Espinoza & Jones, 2021; Jones & Espinoza, 2022), emcee (Foreman-Mackey et al., 2013), matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), PandExo (Batalha et al., 2017), SciPy (Virtanen et al., 2020)
[…]
October 10, 2024
@@ 1 Introduction
Atmospheres are not perfectly uniform. Observational surveys of brown dwarfs and planetary mass objects have found robust evidence that their atmospheres are heterogeneous and variable (e.g., Buenzli et al., 2014; Metchev et al., 2015; Apai et al., 2017). For transiting exoplanets specifically, there is abundant evidence for day-to-night variation in atmospheric temperature and cloudiness (e.g., Beatty et al., 2019; May et al., 2022, among many others) driven by the contrast in irradiation received by the day and night hemispheres. Three-dimensional atmospheric circulation models also predict that this day-to-night contrast will drive differences between the morning and evening terminator regions (e.g., Kataria et al., 2016; Line & Parmentier, 2016; Powell et al., 2019), which are known as “limb asymmetries” because these terminators form the limbs of the planetary disk when seen in transit.
@@
The James Webb Space Telescope (JWST) has proven capable of directly detecting limb asymmetry at low spectral resolution as well. The observational signature of limb asymmetry in JWST transit observations are apparent time of conjunction variations with wavelength (Feinstein et al., 2022; Rustamkulov et al., 2022) and changes in the chromatic transit depth between ingress and egress (Murphy et al., 2024; Espinoza et al., 2024; Delisle & et al., 2024, in review). One advantage of JWST is that it captures the transmission spectra, both overall and of each limb’s contribution, over a much wider wavelength range than is generally possible in high-resolution spectroscopy from the ground. This enables more detailed retrievals of the atmospheric temperature–pressure structure, molecular abundances, and aerosol properties in each spatial region using JWST. Extracting these separate contributions is vital to studying and understanding the circulation of heat and molecules within atmospheres. Also, previous analyses have shown that, if such differences are ignored, the properties retrieved from observed data can be biased significantly from the truth (Feng et al., 2016; Caldas et al., 2019; Taylor et al., 2020).
@@ 2.1 Defining the Problem
Consider an example planet with asymmetric limbs, where each limb is represented by a semicircle joined to the other along its flat edge (as in Figure 1, inspired by Espinoza & Jones, 2021; Jones & Espinoza, 2022). The evening limb has radius $R_{\mathrm{evening}}$ and the morning limb radius differs by some amount $\Delta R$, so that $R_{\mathrm{morning}}=R_{\mathrm{evening}}+\Delta R$. This $\Delta R$ can be positive or negative depending on which limb is larger, and herein we will assume $\Delta R$ is positive, so that the example planet’s morning limb is larger. While this construction simplifies the geometry, we note it is not a strictly realistic model, as it creates a discontinuity in the atmosphere at the planet’s poles that grows with $\Delta R$. However, this simplified construction is likely adequate given current data quality, which we discuss later.
An observer seeking to study this planet’s limb asymmetry must measure both $R_{\mathrm{evening}}$ and $R_{\mathrm{morning}}$ [^1] In practice, it is really the corresponding planet–star radius ratios $R_{\mathrm{evening}}/R_{\star}$ and $R_{\mathrm{morning}}/R_{\star}$ that are measured from a transit light curve. as a function of wavelength, thus constructing separate transmission spectra for each limb. We assume that previous observers of this planet were not aware of – or not capable of – measuring limb asymmetry on the planet, and fit their data assuming the planet had uniform limbs. The planetary radius that these analyses will have inferred is the radius of a uniform circle with the same occulting area – the same transit depth – as the asymmetric-limb disk, which we call the “effective radius” $R_{\mathrm{eff}}$, calculated as
@@
Figure 1: Exaggerated schematic illustrating the difference in the transit light curves of an asymmetric-limb planet and a uniform-limb planet. Limb darkening is omitted for visual simplicity. Both planets share the same time of conjunction. Due to the difference in their leading and trailing limb radii though, the transit contact times of each planet differ. An observer that measures the light curve of the asymmetric-limb planet, which is shown in blue, and fits their data using a uniform-limb transit model would infer a biased time of conjunction because these changes in the contact times effectively shift the transit earlier or later in time. This is the origin of the degeneracy between limb asymmetry and variations or uncertainty in a planet’s time of conjunction.
To best intuit the degeneracy between $\Delta R$ and $\Delta t_{c}$, it is useful to look at the light curve of the asymmetric-limb planet compared to its uniform-limb counterpart. Figure 1 shows an exaggerated example of each at the bottom, which are aligned in time with the sequence of disk schematics at the top. In this example, both planets have the exact same true time of conjunction and orbital speed, and are moving from left to right across the stellar disk. The asymmetric-limb planet will begin to transit before the uniform-limb planet because its leading limb, which is $R_{\mathrm{morning}}$, is larger than $R_{\mathrm{eff}}$. In other words, because it has a larger limb, the asymmetric-limb disk makes first contact (T1) with the stellar disk before the uniform-limb disk does. Similarly, since the asymmetric-limb disks’ trailing edge, $R_{\mathrm{evening}}$, is smaller than $R_{\mathrm{eff}}$, it is still “ahead” of the trailing edge of the uniform-limb disk, so it will also contact the stellar disk first during the end of ingress (T2). The exact same sequence of events occurs during transit egress for the third (T3) and fourth (T4) contact points. Also note that the shape of the ingress and egress in each light curve is slightly different due to the different shape of each disk, but they both reach the same depth during full transit since both disks occult the same total area.
@@ 2.2 Times of First and Last Contact
We assume that each planet rotates about a single axis that is oriented parallel to its orbital axis and lies in the plane of the page. The following derivation will hold reasonably well even if there is small rotation–orbital axis misalignment, though such misalignment is unlikely for short-period exoplanets, due to tidal interactions with the host star (e.g. Hut, 1981; Correia & Laskar, 2010). As shown in Figure 1, we construct each planetary disk as two semicircles whose straight edges are joined along the planet’s rotation axis. Here, we explicitly define the “center” of each planet as the point where the rotation axis meets the two lines which bisect the arcs of each semicircle, as indicated in Figure 1. Under this construction, these bisectors are parallel to the planets’ transit trajectories. To aid our derivation, we will work in a Cartesian coordinate system as is also shown in Figure 1. The x-axis is aligned along the planet’s transit chord, so that the center of the planet is always moving directly along the x-axis.
The point of conjunction is the position of the center of each planet at its time of conjunction, defined as when the center of the planet is closest to the center of the stellar disk. We define this point of conjunction to be the origin of our coordinate system, so that the y-axis passes through this point and is orthogonal to both planets’ transit trajectories. Note that, in Figure 1, we assume both planets transit directly along the stellar equator, but this need not be true in general. Our coordinate system is invariant to general changes in the planets’ orbital orientations, as the x-axis can be translated vertically along the stellar disk and the y-axis rotated such that the x-axis is always horizontal.
@@
Figure 2: Projected geometry of a planet with asymmetric limbs at transit first contact. A line can be drawn from the center of the star to the center of the planet that passes through the point of first contact. This line defines the hypotenuse of a triangle that we use to derive how the time of first contact changes as the leading limb radius changes in Equation 2.2.
When the asymmetric-limb planet makes first contact, assuming it has a larger leading limb, the uniform-limb planet will still need to travel some extra projected distance in the $x$-direction before it too makes first contact. Let us hold the asymmetric-limb planet fixed at first contact, and let the uniform-limb planet move forward in time until it too makes first contact. Then, we solve for the difference in these times via this extra projected distance $\Delta x$ that the uniform-limb planet travels. When the planets are transiting along the stellar equator, as is shown in Figure 1, this is simply $\Delta x=R_{\mathrm{morning}}-R_{\mathrm{eff}}\approx\Delta R/2$. For nonzero impact parameters, estimating $\Delta x$ is more involved. Figure 2 shows the projected geometry at the point of first contact for an asymmetric-limb planet with nonzero impact parameter. We can draw a triangle with a hypotenuse from the center of the stellar disk, through the point of first contact, to the center of the planetary disk. The length of this hypotenuse is $R_{\star}+R_{\mathrm{morning}}$ and it makes an angle $\theta_{1}$ with respect to the stellar equator. The projected vertical distance between the center of the planet and the stellar equator is $bR_{\star}$, where $b$ is the impact parameter. Let us define the position of the center of the asymmetric-limb planet when its leading limb makes first contact as $x_{a}$, and likewise for the uniform-limb planet as $x_{u}$. The extra distance traveled by the uniform-limb planet is {dmath} Δx = x_a - x_u
@@
Here, $R_{p,i}$ can be $R_{\mathrm{morning}}$ or $R_{\mathrm{eff}}$. We can convert Equation 3 to the cosine of the angle
@@
The same derivation can be done for the time of fourth contact by swapping the morning limb radius for the evening limb radius. In our construction, the uniform-limb planet will also make fourth contact after the asymmetric-limb planet, now because the uniform-limb radius is larger than the evening-limb radius. Changing out terms, we then have {dmath} ΔT_4 = P2 πa ( (R_⋆+ R_eff )^2 - (b R_⋆)^2 - (R_⋆+ R_evening )^2 - (b R_⋆)^2 ). By construction, $\Delta T_{1}$ and $\Delta T_{4}$ will always have the same sign. This is because the equivalent-depth uniform radius will always be between the two limb radii, so T1 and T4 will always be earlier or later together. If the asymmetric-limb planet’s morning limb is larger than its evening limb, as we have been using in our constructed example, then $\Delta T_{1}$ and $\Delta T_{4}$ will be negative. If the evening limb is larger, they will be positive. Then, because these contact times are equidistant about the time of conjunction for the uniform-limb planet: {dmath} Δt_c = 12 ( ΔT_1 + ΔT_4 ).
@@ 3.1 Setup
We first simulated a “perfect” transit observation of the exoplanet WASP-39b in order to directly compare to our analytical formulae. We generated a noiseless light curve with data points at a cadence of 0.5 s. For this “perfect” observation, we finely sampled $N$ in 20 evenly spaced steps between 0 and 5.
@@
Parameter (units) WASP-80b WASP-96b WASP-39b WASP-107b
P (days) 3.06785271(19)aafootnotemark: 3.4252565(8)aafootnotemark: 4.0552941(34)fffootnotemark: 5.721488(3)aafootnotemark:
@@
$i$ (degrees) 89.02 $\pm$ 0.11bbfootnotemark: 85.6 $\pm$ 0.2ccfootnotemark: 87.32 $\pm$ 0.17fffootnotemark: 89.7 $\pm$ 0.2jjfootnotemark:
@@
R (R) 0.605 $\pm$ 0.048**footnotemark: 1.15 $\pm$ 0.03**footnotemark: 0.92 $\pm$ 0.03**footnotemark: 0.73 $\pm$ 0.15**footnotemark:
@@
Teq (K) 825 $\pm$ 19bbfootnotemark: 1285 $\pm$ 40ccfootnotemark: 1166 $\pm$ 14fffootnotemark: 770 $\pm$ 60jjfootnotemark:
T14 (hr) 2.131 $\pm$ 0.003bbfootnotemark: 2.4264 $\pm$ 0.0264ccfootnotemark: 2.8032 $\pm$ 0.0192iifootnotemark: 2.753 $\pm$ 0.007jjfootnotemark:
H (km) 226 $\pm$ 44 669 $\pm$ 97 1042 $\pm$ 171 1221 $\pm$ 174
Cadence (s) 14.8 72.0 79.5 20.2
$\sigma$ (ppm) 200 150 200 200
@@
We then simulated more realistic transit observations of each planet using the optimal integration cadences computed using PandExo (Batalha et al., 2017) for a JWST/NIRCam F322W2 observation, and light-curve scatters of 200 ppm. For these simulations, we used a coarser sampling for $N$ of 10 steps between 0 and 5. The cadences are given in Table 1 and ranged from $\sim$15 - 80 s, consistent with those used by real observations of these targets (Bell et al., 2023; Ahrer et al., 2023; Murphy et al., 2024). This light-curve scatter is slightly optimistic for JWST/NIRCam F322W2, as real observations of these targets have achieved $\sim$230-500 ppm scatter (Bell et al., 2023; Carter et al., 2024; Murphy et al., 2024), but it has been achieved by other JWST instruments like NIRISS SOSS (Radica et al., 2023). For WASP-96b, its nominal cadence of 144 s and 200 ppm light-curve scatter led to large noise in our simulation results. We therefore used a 2$\times$ faster cadence of 72 s and slightly smaller scatter of 150 ppm just for WASP-96b, which did not change the underlying result but made it easier to visually present and compare to our analytical results. In all cases, we assumed having an equal time out-of-transit baseline to the published transit duration.
@@
After generating the simulated observed data, we fit these data using a uniform-limb transit model generated by the transit modeling package batman (Kreidberg, 2015), also with uniform limb darkening. We left all parameters fixed to their true values except for the time of conjunction, which was allowed to vary freely. We set the planetary radius to the value that would have the same planetary disk occulting area as the corresponding asymmetric-limb disk, using Equation 1. In the “perfect” noiseless case, we simply fit the data through a least-squares regression. Then in the other cases, we fit the data using the Markov Chain Monte Carlo method using emcee (Foreman-Mackey et al., 2013), sampling just the time of conjunction for 5500 steps which was more than sufficient for each run’s sampling to converge.
@@
We chose to use uniform limb darkening when generating the reference light curves, meaning the stellar disks are completely uniform in brightness. This choice was partially motivated by computational efficiency, as the overhead initializations of catwoman and batman are significantly slower when limb darkening must be computed. It was also motivated by our desire to treat each planet uniformly, as different host stars exhibit different limb darkening that may yield different results depending on the assumed cadence or noise of the light curve. We did test various limb darkening laws, including quadratic and logarithmic, and various sets of limb darkening coefficient values. We found that the precise choice of limb darkening did not change our result so long as it was well known, because we set the limb darkening when generating the reference light curve and kept it fixed for all later steps, thus assuming perfect knowledge of limb darkening during the light-curve fitting. Uncertainty in the limb darkening coefficients is of concern to real observations, but is beyond the scope of our analysis and can be mitigated using stellar models in practice.